OpenStudy (anonymous):

A certain field is a rectangle with a perimeter of 1040 feet. The length is 180 feet more than the width. find the width and length of the rectangular field.

OpenStudy (anonymous):

whats the equation of a rectangle's perimeter? do you know it?

OpenStudy (anonymous):

is this a multiple choice question?

OpenStudy (anonymous):

P=2(L+W)

OpenStudy (anonymous):

L=(180+w),L is the length

OpenStudy (anonymous):

P=1040

OpenStudy (anonymous):

1040=2[w+(180+w)] 1040/2=w+180+w 520=2w+180 520-180=2w 340=2w 340/2=w w=170

OpenStudy (anonymous):

L=w+180 L=170+180 L=350

OpenStudy (anonymous):

L=w+180 L=170+180 L=350